what is the next 2 terms?
9, 3, 1, 1/3, __, __

Answer by thecampusnerd(13)

Oh I love these things!

Ok, if you'll notice, you're dividing each term by 3.

9/3 = 3, 3/3 = 1, 1/3 = 1/3...

So, you'll need to divide 1/3 by 3. That's (1/3) / (3/1) which equals (1/3) x (1/3) because you have to invert the fractions and change the operation sign to its opposite. Now multiply straight across. Your answer is 1/9. That's also your next term.

Now divide 1/9 by 3. Do the same thing I just did previously, except change the numbers. Your answer is 1/27.